A preliminary look at the result of the HVAC mitigation actions at NEB and WEB. We measure a reduction of acoustic and also (a bit) seismic noise in the experimental halls of both NEB and WEB.
A reduction is clearly seen between 20 and 200Hz.
NEB is shown in Figure 1 and 2 (zoom), WEB is shown in Figure 3 and 4 (zoom).
NOTE: in all the plots, blue curves are before the mitigation, PURPLE curves are AFTER the mitigation.
Consequences of the intervention on INF signals, previously measured by sensors installed in the ducts removed (see plot1).
LEGENDA:
All the sensors are installed on the UTA outlet duct (PRES_OUT: differential pressure between UTA and external environment, TE_OUT, TE2_OUT: air temperature, VEL_OUT: air velocity)
WEB:
Drop of INF_WEB_PRES_OUT: not completely recovered after my intervention (yesterday afternoon), the pression could be lower, perhaps, to be further checked.
Spikes on INF_WEB_TE_OUT: discovered and repaired this morning.
Relevant differences on INF_WEB_TE2_OUT and INF_WEB_VEL_OUT: the sensor is really installed in a very unusual position for a temperature and air velocity sensor!!! (see pict1)
NEB:
Drop of INF_NEB_PRES_OUT: not recovered (sensor broken, to be replaced).
INF_WEB_TE_OUT: one cable was unscrewed from it's connector, recovered few minutes ago.
INF_WEB_TE2_OUT: the temperature has large fluctuation respect to the past, to be compared with INF_WEB_TE_OUT.
Drop of INF_NEB_VEL_OUT: recovered in two times. This sensor must be placed perpendicular to the air flow and it was slightly rotated (completely recovered - with the help of a stair - this morning).